Monday November 28
The day shift began with the Tevatron (TeV) sector B17 work, and with MiniBooNE, NuMI, and the Meson MTest and MCenter experiments getting no beam due to Linac tuning.
Operators accessed the Booster and discovered a LCW leak and repairs began. Booster experts also worked on a kicker (MKS05), a power supply (GMPS), and replaced a RF station (BRF6). MiniBooNE experts worked on their detector. Booster was turned back on by late afternoon, followed by beam to NuMI and MiniBooNE and then later beam to MCenter. The Antiproton Source (Pbar) and the Recycler conducted studies all evening.
Tuesday November 29
The midnight shift began with TeV B17 work, with Pbar and Recycler conducting studies, and with MiniBooNE, NuMI, and MCenter taking beam.
Operations shut down beam at 6 AM for beam line accesses to Linac, Booster, Main Injector (MI), Pbar, and the TeV F-Sector and transfer hall. During the day shift, MI-40 enclosure was found with four inches of water on the floor, the Meson MS2 service building lost its power, and technicians repaired the P2/F sector transfer line vacuum problems. Pbar experts reported on the evening shift that they had repaired the Lithium Lens.
Wednesday November 30
The midnight shift began with operations beginning turn on for all machines but the TeV.
Operation established beam to MiniBooNE, NuMI, and Switchyard (SY) around 4 AM.
Pbar began a reverse proton tune up at 4:26 AM.
A vacuum expert reported at 7:41 AM that there's still a vacuum leak in Booster.
Operations put Pbar into stacking mode at 7:54 AM, but this was at low intensity and an expert was tuning.
The Plans for Wednesday
The plans for the present are to get the machines back to normal operations as soon as possible.
Accelerator Update Archive
For Tevatron luminosity charts and the current status of Fermilab's
accelerators and detectors (live!), please go to Fermilab Now.
Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: email@example.com.